OpenStudy (anonymous):

\[\int\limits_{1}^{\infty} \left( \frac{\ln(x)}{x} \right)^{2011}dx\]

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

myininaya (myininaya):

hey is this like a, whats it called, published problems or something like that? like i see 2011 i done a 2007 problem in 2007 lol

OpenStudy (anonymous):

oh haha its from a math tournament in 2011

OpenStudy (anonymous):

probly why they chose that power

myininaya (myininaya):

i should show you the problem i did in my senior seminar class

OpenStudy (anonymous):

id like to see it if you have it

myininaya (myininaya):

\[\lim_{n \rightarrow \infty}(1^\frac{1}{n}+2^\frac{1}{n}+...+2006^\frac{1}{n}-2007)\]

OpenStudy (anonymous):

no power on 2007 right?

myininaya (myininaya):

oops to the nth power

myininaya (myininaya):

the whole thing is to the n powee 2007 is the 1st power

myininaya (myininaya):

\[\lim_{n \rightarrow \infty}(1^\frac{1}{n}+2^\frac{1}{n}+...+2006^\frac{1}{n}-2007)^n\]

OpenStudy (anonymous):

well i think its going to take some thought

OpenStudy (anonymous):

im trying some e^ln(...) things

myininaya (myininaya):

i think yours is going to take some thought as well

OpenStudy (anonymous):

woops just broke maple. I tried to define the function in the limit using a for loop

OpenStudy (anonymous):

the integral i posted doesnt work in mathematics software either

OpenStudy (anonymous):

how long did it take you to solve this problem?

myininaya (myininaya):

a semester lol

myininaya (myininaya):

i didn't really solve it. i had help from the instructor but i understand it now

OpenStudy (anonymous):

yikes well then let me know how to start it haha

myininaya (myininaya):

you will need to use or i used Let T(x)=a^x T'(0)=lim x->0 (f(x)-f(0))/(x-0)=(a^x-1)/x=lna. this should really help you get started i think.

myininaya (myininaya):

let x=1/n then n goes to infinity

myininaya (myininaya):

i forgot to type x->0 on that one part if you didn't notice lim x->0 (a^x-1)/x=lna

myininaya (myininaya):

have you actually done the problem you typed above?

OpenStudy (anonymous):

yeah i have

myininaya (myininaya):

how long did it take you

OpenStudy (anonymous):

about an hour of doing wrong things and then I did something that worked.

myininaya (myininaya):

lol

OpenStudy (anonymous):

do you want the general method to use?

myininaya (myininaya):

let me look at it for a few seconds i know how to integrate lnx/x using substition something like this looks like we have to use integration by parts it will be long but i could probably write it as a summation after figuring out the pattern and then figure out the limit of the summation

myininaya (myininaya):

i have wrote it as int(u^2011*e^(-2010u) du)

OpenStudy (anonymous):

i did that too I couldn't get much further that way though

myininaya (myininaya):

there is a better way?

OpenStudy (anonymous):

yes substitution doesn't work, integration by parts works out very well

myininaya (myininaya):

skip the substitution and go straight for the integration by parts?

myininaya (myininaya):

instead of using both?

OpenStudy (anonymous):

yeah just int by parts, and if you pick the correct u, dv there's a really nice pattern so you dont have to do it 2011 times

myininaya (myininaya):

lol yes 2011 is basically infinity (countable)

OpenStudy (anonymous):

haha yeah it would be no fun, aka maple gave up after 5 minutes

myininaya (myininaya):

let maple run all night and it didn't ever give me a limit for that 2007 problem

OpenStudy (anonymous):

haha computers arent that smart yet

OpenStudy (anonymous):

theyre pretty smart though

OpenStudy (anonymous):

ok well I should get to bed, talk to you later

myininaya (myininaya):

me too someone is waiting on me lol

OpenStudy (anonymous):

good night :)

myininaya (myininaya):

gn

myininaya (myininaya):

i got -2011!*e/(2010^2010)

myininaya (myininaya):

i think this is right

myininaya (myininaya):

oops its divergent sorry

myininaya (myininaya):

i wrote it from 0 to infinity instead of 1 to infinity so we just have to look at lnb as b goes to infinity lnb->infinity

OpenStudy (anonymous):

\[u=\ln(x)^n \rightarrow du=\frac{n*\ln(x)^{n-1}}{x}\] \[dv=\frac{1}{x^n} \rightarrow v=-\frac{1}{(n-1)x^{n-1}}\] so our integral is: \[\frac{\ln(x)^n}{(n-1)x^{n-1}}+\frac{n}{n-1}\int\limits_{1}^{\infty}\frac{\ln(x)^{n-1}}{x^n}\] The first term goes to zero. the second interal we can perform the same integration by parts until we get: \[\frac{n!}{(n-1)^n}\int\limits_{1}^{\infty}\frac{1}{x^n}\] this integral is 1/(n-1) so finally: \[\int\limits_{1}^{\infty}\left( \frac{\ln(x)}{x} \right)^n=\frac{n!}{(n-1)^{n+1}}=\frac{2011!}{2010^{2012}}\] the general version is true when n>0

OpenStudy (anonymous):

sorry, when n>1